# Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery

### v3

@@ A.2 Copper Latent Factor Correlations
Table 7: Pearson Correlation Coefficients: Copper Latent Factors vs. Selected Features (Test Set).
@@
Global Mfg. PMI (Change) $0.75$ $-0.10$ $0.05$ $-0.20$
LME Inventories (Change) $-0.65$ $0.05$ $0.15$ $0.10$
USD Index (DXY) (Change) $-0.25$ $0.70$ $-0.10$ $0.30$
Supply Disruption Index (Proxy) $0.10$ $0.00$ $-0.80$ $0.15$
VIX (Market Volatility) $-0.15$ $0.35$ $0.55$ $0.65$
@@
  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with a (hypothetical) Supply Disruption Index (where lower values of the index indicate more severe disruptions). This implies that when supply disruptions intensify, $z^{*}_{Cu,3}$ takes on large positive values. Its moderate positive correlation with VIX (+0.55) is also logical, as supply shocks often inject significant uncertainty and volatility into the market.
@@
A.3 Crude Oil Latent Factor Correlations
@@
Table 8: Pearson Correlation Coefficients: Crude Oil Latent Factors vs. Selected Features (Test Set).
@@
EIA US Crude Inv. (Change) $-0.70$ $0.10$ $0.05$ $-0.15$
US Gasoline Demand (Proxy Change) $0.60$ $0.05$ $0.10$ $0.20$
Geopol./OPEC+ Event (Proxy) $0.15$ $0.80$ $0.30$ $0.00$
OVX (Oil Volatility) $-0.10$ $0.45$ $0.75$ $-0.25$
Global Growth Lead Ind. (Change) $0.20$ $0.10$ $-0.15$ $0.70$
@@
  • Factor $z^{*}_{Oil,1}$ (US Demand, Inventory, and Refining Dynamics): This factor shows a strong negative correlation (-0.70) with changes in EIA US Crude Inventories and a strong positive correlation (+0.60) with a proxy for US Gasoline Demand. This clearly links $z^{*}_{Oil,1}$ to the state of the US oil market balance: higher values of this factor correspond to falling inventories and/or rising demand, typically bullish for WTI prices.
  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with a proxy for Geopolitical/OPEC+ supply-tightening events (where a positive value for the proxy indicates an event expected to reduce supply or increase risk premium). It also correlates moderately (+0.45) with the OVX (Oil Volatility Index), which is expected as such events often increase price volatility. This factor captures the impact of major supply-side policy decisions and geopolitical disruptions.
@@
  • Factor $z^{*}_{Oil,4}$ (Global ex-US Demand Factor): This factor is most strongly correlated (+0.70) with changes in a Global Economic Growth Leading Indicator. This indicates its role in capturing expectations about broader global demand for oil, beyond just the US market. Its moderate positive correlation with the US Gasoline Demand proxy (+0.20) is also consistent with general demand trends.
These interpretations, based on correlations, provide a plausible understanding of how the RSAE model might be decomposing market information into sparse, meaningful drivers. Further validation using more formal attribution methods (as discussed in Section 6.3) would be required to confirm these relationships more rigorously.

### v4

@@ C.2 Copper Latent Factor Correlations
Table 15: Pearson Correlation Coefficients: Copper Latent Factors vs. Selected Features (Test Set). Units: correlation coefficients (unitless).
@@
Global Mfg. PMI (Change) $0.75$ 0.10-0.10 $0.05$ 0.20-0.20
LME Inventories (Change) 0.65-0.65 $0.05$ $0.15$ $0.10$
USD Index (DXY) (Change) 0.25-0.25 $0.70$ 0.10-0.10 $0.30$
Supply Disruption Index (Observed) $0.10$ $0.00$ 0.80-0.80 $0.15$
VIX (Market Volatility) 0.15-0.15 $0.35$ $0.55$ $0.65$
@@
  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with the observed Supply Disruption Index (Appendix C), where lower values denote more severe outages. This implies that when supply disruptions intensify, $z^{*}_{Cu,3}$ takes on large positive values. Its moderate positive correlation with VIX (+0.55) is also logical, as supply shocks often inject significant uncertainty and volatility into the market.
@@
C.3 Crude Oil Latent Factor Correlations
@@
Table 16: Pearson Correlation Coefficients: Crude Oil Latent Factors vs. Selected Features (Test Set). Units: correlation coefficients (unitless).
@@
EIA US Crude Inv. (Change) 0.70-0.70 $0.10$ $0.05$ 0.15-0.15
US Gasoline Demand (DOE product supplied) $0.60$ $0.05$ $0.10$ $0.20$
Geopol./OPEC+ Event Index $0.15$ $0.80$ $0.30$ $0.00$
OVX (Oil Volatility) 0.10-0.10 $0.45$ $0.75$ 0.25-0.25
Global Growth Lead Ind. (Change) $0.20$ $0.10$ 0.15-0.15 $0.70$
@@
  • Factor $z^{*}_{Oil,1}$ (US Demand, Inventory, and Refining Dynamics): This factor shows a strong negative correlation (-0.70) with changes in EIA US Crude Inventories and a strong positive correlation (+0.60) with DOE product supplied (gasoline demand). This clearly links $z^{*}_{Oil,1}$ to the state of the US oil market balance: higher values correspond to falling inventories and/or rising demand, typically bullish for WTI prices.
  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with the Geopolitical/OPEC+ event index, where positive values signal supply-tightening announcements. It also correlates moderately (+0.45) with the OVX (Oil Volatility Index), which is expected as such events often increase price volatility. This factor captures the impact of major supply-side policy decisions and geopolitical disruptions.
@@
  • Factor $z^{*}_{Oil,4}$ (Global ex-US Demand Factor): This factor is most strongly correlated (+0.70) with changes in a Global Economic Growth Leading Indicator. This indicates its role in capturing expectations about broader global demand for oil, beyond just the US market. Its moderate positive correlation with the DOE gasoline product supplied series (+0.20) is also consistent with general demand trends.
These interpretations, based on correlations, provide a plausible understanding of how the SLFF model might be decomposing market information into sparse, meaningful drivers. Further validation using more formal attribution methods (as discussed in Section 5.5) would be required to confirm these relationships more rigorously.
Appendix D Event Studies and Counterfactuals
This appendix lists the event windows (pre/post days), factor responses, and forecast impacts used in the counterfactual study. Each table includes the event definition (e.g., OPEC+ surprise cuts, ICSG mine shutdown announcements, CFTC positioning extremes), the latent factors most affected, the t-statistics of the factor change relative to a randomization null, and the induced shift in 5-day and 22-day forecasts. Scripts to regenerate these tables are part of the released interpretability notebooks.